OpenStudy (anonymous):

Find tangent line to f(x) = x + 2 sin x at x =  6 and x =  3 .

OpenStudy (anonymous):

its suppose to be pi/6 and pi/3

OpenStudy (anonymous):

slope = f'(x)=1+2cosx=1+2cos(pi/6)=1+sqrt(3) at x=pi/6 f(x)=1+2sinpi/6=1+1=2 eq is y-2=(1+sqrt3)(x-pi/6)

OpenStudy (anonymous):

for x=pi/3, similar method find slope of the line and use slope point form of the line

OpenStudy (anonymous):

thank you